OpenStudy (anonymous):

Owen scored 92 points on a 100-point test and 27 on a 30-point quiz. On which did he get a larger fraction of the total points correct? Show your work

OpenStudy (anonymous):

Change them to fractions: \(\Huge\frac{92}{100}\) \(\Huge\frac{27}{30}\) Can you divide 92 \(\div\) 100 and 27 \(\div\) 30?

OpenStudy (anonymous):

i love you.

OpenStudy (anonymous):

92 \(\div\) 100 = 0.92 27 \(\div\) 30 = 0.90 0.92 = \(\huge\frac{92}{100}\) 0.90 = \(\huge\frac{90}{100}\) Which one is bigger?

OpenStudy (anonymous):

92/10

OpenStudy (anonymous):

just needed to get my answer approved

OpenStudy (anonymous):

Right, so "92 points on a 100 question test" has a higher grade than "27 points on a 30 question test". You can give meals by clicking 'Best Response'.

OpenStudy (anonymous):

\(\huge\cal\rlap{\color{lime}-\color{lime}G\color{lime}r\color{lime}e\color{lime}e\color{lime }n\color{lime}ツ}{\color{black}{\;-Green~ツ}}\)

OpenStudy (anonymous):

i wanna eat yaa so lemme eat ya

OpenStudy (anonymous):

<:

OpenStudy (anonymous):